Numerical Analysis I, Fall 2020 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 04

Goal: Study both theoretical and practical aspects of Newton’s method and its extensions. 1. Section 2.3: Problems 13(b,c).